OpenStudy (anonymous):

help me solve this 6x+8y=18 and 5x-8y=-34

OpenStudy (kinggeorge):

Since you have 8y in the first, and -8y in the second, simply add the two equations together. You get 11x = -16 or x = -16/11. Then substitute x back in, and solve for y.

OpenStudy (gw2011):

6x+8y=18 5x-8y=-34 Add the equations to get: 11x=-16 x=-16/11 Substitute the x value above into one of the equations to get: 6x+8y=18 (6)(-16/11)+8y=18 -96/11+8y=18 8y=18+96/11 8y=198/11+96/11 8y=294/11 Divide both sides by 8 to find y: y=294/88